OpenStudy (anonymous):

Use integration, direct comparison test, or the limit comparison test to test the integral for convergence. integral sin x/(sqrt. pi - x) dx ; int. 0 to pi

OpenStudy (anonymous):

\[\int\frac{sin(x)}{\sqrt{\pi-x}}dx\]